# Indic-TunedLens: Interpreting Multilingual Models in Indian Languages

### v1

@@ Abstract
Multilingual large language models (LLMs) are increasingly deployed in linguistically diverse regions like India, yet most interpretability tools remain tailored to English. Prior work reveals that LLMs often operate in English centric representation spaces, making cross lingual interpretability a pressing concern. We introduce Indic-TunedLens, a novel interpretability framework specifically for Indian languages that learns shared affine transformations. Unlike the standard Logit Lens, which directly decodes intermediate activations, Indic-TunedLens adjusts hidden states for each target language, aligning them with the target output distributions to enable more faithful decoding of model representations. We evaluate our framework on 10 Indian languages using the MMLU benchmark and find that it significantly improves over SOTA interpretability methods, especially for morphologically rich, low resource languages. Our results provide crucial insights into the layer-wise semantic encoding of multilingual transformers. Our model is available at https://huggingface.co/spaces/AnonymousAccountACL/IndicTunedLens. Our code is available at https://github.com/AnonymousAccountACL/IndicTunedLens.
@@
Mihir Panchal1, Deeksha Varshney2, Mamta3, Asif Ekbal2 1 Dwarkadas Jivanlal Sanghvi College of Engineering, Mumbai, India 2 Indian Institute of Technology Jodhpur, Jodhpur, India 3 King’s College London, London, UK mihirpanchal5400@gmail.com, deeksha@iitj.ac.in, mamta.name@kcl.ac.uk, asif@iitj.ac.in
@@ 4 Experimental Setup and Analysis
Indic-TunedLens is trained using the Sarvam-1 model as the base model. All the implementations are carried out in Pytorch  [^2] https://pytorch.org/. To train and test Indic-TunedLens, we utilize two datasets that provide comprehensive multilingual coverage across Indian languages. We train our model on eleven languages, Bengali, English, Gujarati, Hindi, Kannada, Malayalam, Marathi, Panjabi, Nepali, Tamil and Telugu from the Sangraha dataset (Rahman Khan et al., 2024). For evaluation, we employ a curated subset of the multilingual Massive Multitask Language Understanding (MMLU) dataset (Dac Lai et al., 2023) adapted for Indian languages, comprising Bengali, English, Gujarati, Hindi, Kannada, Malayalam, Marathi, Nepali, Tamil and Telugu samples. The inclusion of Panjabi as a train only language enables assessment of cross lingual transfer capabilities to unseen but related languages. Detailed dataset statistics and implementation details are provided in Appendices B.1 and B.2, respectively.
@@ 5 Results
To assess the effectiveness of Indic-TunedLens for multilingual transformer interpretability, we conduct a comprehensive comparative analysis against the standard Logit Lens approach across 10 Indian languages (detailed in Appendix B).
Figure 1: Entropy heatmap showing inconsistent patterns and high variance across layers for standard Tuned Lens, revealing fragmented attention patterns and poor alignment with Indian language processing pathways
Figure 2: Entropy heatmap showing entropy coherent reduction and information consolidation across layers in Indic-TunedLens, reflecting improved semantic capture for Indian languages.
Layer-wise uncertainty reveals language-specific meaning formation The standard Tuned Lens approach, originally designed for English centric models, demonstrates significant limitations when applied to Indian language processing, as illustrated in Figure 1. We show an example question (Figure 3) about monopoly concentration ratios from the Hindi MMLU dataset [^3] https://huggingface.co/datasets/alexandrainst/m_mmlu/ viewer/hi. The entropy (calculated as discussed in Section 4) heatmap which measures uncertainty in the model’s predicted token distribution at each layer reveals inconsistent patterns with high variance across layers, indicating poor alignment between the tuning mechanism and the model’s internal processing pathways for Indian languages. The visualization shows fragmented attention patterns, particularly in the middle layers (8-16), where semantic consolidation should occur, and scattered high entropy regions without clear progressive information refinement. This degraded performance stems from a projection mismatch: while the model internally encodes morphologically rich and semantically meaningful representations for Indian languages, the English-centric Tuned Lens fails to align these representations with the output vocabulary space.
@@
Maximum improvements occur in early layers (1-8) with language specific patterns reflecting morphological complexity. The heatmap in Figure 4 shows that Bengali, Nepali, and Hindi exhibit the strongest improvements in layers 1-4 (0.04-0.06 accuracy gain), corresponding to the morphological analysis phase where rich inflectional systems are processed. English shows concentrated improvements at layer 1 (0.06 gain) and more modest gains throughout middle layers, reflecting its analytic morphology. Gujarati and Telugu demonstrate sustained improvements across layers 1-10 (0.02-0.04 gain), while Malayalam and Kannada show more distributed improvement patterns extending into middle layers (5-15). These layer specific improvement patterns enable linguistic hypotheses about where different linguistic phenomena are computed, suggesting that morphological analysis and early semantic composition occur in early layers, while deeper layers support more complex compositional and syntactic integration. This process happens earlier in languages like Telugu and Tamil, which form words by joining many meaningful parts together, compared to languages with more isolated word structures. The variation in improvement patterns across languages suggests that the learned transformations adapt to language specific processing dynamics, capturing the distinct computational trajectories by which different Indic languages resolve ambiguity and build compositional meaning. This differential effectiveness across layers provides insights into the linguistic mechanisms that the base model employs for each language family, with Dravidian languages showing deeper processing requirements than Indo-Aryan languages. These findings demonstrate that interpretability methods cannot be language-agnostic: effective analysis of multilingual LLMs requires language-family-aware lenses that respect typological differences in morphological and compositional structure.
@@
Indic-TunedLens demonstrates superior accuracy over Logit Lens across all Indian languages, with particularly pronounced improvements in early and middle layers. As shown in Figure 5, the standard Logit Lens exhibits minimal accuracy in early layers (0-15) across all languages, with values remaining below 0.02 for most Indian languages. The accuracy only begins to increase substantially after layer 20, reaching peak performance at layer 27 before dropping sharply at the final layer. In contrast, Indic-TunedLens shows immediate interpretability from layer 1, with English achieving 0.13 accuracy, followed by Telugu (0.11) and Gujarati (0.10). The framework maintains consistent accuracy between 0.03-0.06 throughout middle layers (5-20) for Indian languages, demonstrating robust intermediate representation capture. These findings suggest that language specific affine transformations enable meaningful interpretation of hidden states from the earliest layers of processing, whereas English centric approaches fail to capture the morphological richness and syntactic diversity inherent in Indian languages until much deeper in the network. The superior early layer performance of Indic-TunedLens indicates that specialized training on diverse Indic scripts creates representations better aligned with the model’s internal processing pathways for these languages, enabling more accurate decoding of intermediate semantic states. Individual language specific accuracy curves across all layers are provided in the appendix C.1.
@@
Indic-TunedLens consistently positions correct tokens at higher ranks across all layers, indicating improved confidence in predictions. Figure 6 reveals that Logit Lens maintains average ranks between 4-5 across layers 5-25 for most languages, with Malayalam showing particularly poor performance (ranks 4.5-5.8). The ranks only improve dramatically at layer 27, converging to 3.5-4 across languages. Conversely, Indic-TunedLens exhibits more favorable ranking patterns throughout the network, with most languages maintaining average ranks between 3.9-4.7 across middle layers. Notably, Hindi demonstrates superior rank positioning (4.5-5 in early layers, improving to 3.5-4.0 in later layers), while Malayalam shows the most substantial improvement from its Logit Lens baseline, achieving ranks of 4.2-4.8 in layers 10-25. The consistently better ranking performance of Indic-TunedLens across all languages demonstrates that the learned affine transformations more effectively align intermediate representations with the final vocabulary distribution, placing correct tokens among the top candidates earlier in the decoding process. This improved ranking behavior is particularly critical for morphologically rich languages where token level disambiguation requires integration of multiple linguistic features, and suggests that Indic-TunedLens better captures the hierarchical semantic refinement that occurs across transformer layers for Indian languages. Detailed rank distribution plots for each language are available in the appendix C.2.
@@ 7 Conclusion
Appendix A Appendix
Appendix B Experimental Setup
B.1 Dataset
@@ B.2 Implementation Details
Appendix C Detailed Analysis
C.1 Language Wise Layer-wise Accuracy Comparison between Indic-TunedLens and LogitLens
@@
Figure 17: Layer-wise Accuracy Comparison for Telegu
C.2 Language Wise Average Rank of Correct Predictions between Indic-TunedLens and LogitLens
@@
Figure 27: Average Rank of Correct Predictions for Telegu

### v2

@@ Abstract
Multilingual large language models (LLMs) are increasingly deployed in linguistically diverse regions like India, yet most interpretability tools remain tailored to English. Prior work reveals that LLMs often operate in English centric representation spaces, making cross lingual interpretability a pressing concern. We introduce Indic-TunedLens, a novel interpretability framework specifically for Indian languages that learns shared affine transformations. Unlike the standard Logit Lens, which directly decodes intermediate activations, Indic-TunedLens adjusts hidden states for each target language, aligning them with the target output distributions to enable more faithful decoding of model representations. We evaluate our framework on 10 Indian languages using the MMLU benchmark and find that it significantly improves over SOTA interpretability methods, especially for morphologically rich, low resource languages. Our results provide crucial insights into the layer-wise semantic encoding of multilingual transformers. Our model is available at https://huggingface.co/spaces/MihirRajeshPanchal/IndicTunedLens. Our code is available at https://github.com/MihirRajeshPanchal/IndicTunedLens.
@@
Mihir Panchal1, Deeksha Varshney2, Mamta .3, Asif Ekbal4 1 Dwarkadas Jivanlal Sanghvi College of Engineering, Mumbai, India 2 Indian Institute of Technology Jodhpur, Jodhpur, India 3 King’s College London, London, UK 4 Indian Institute of Technology Patna, Patna, India mihirpanchal5400@gmail.com, deeksha@iitj.ac.in, mamta.name@kcl.ac.uk, asif@iitp.ac.in
@@ 4 Experimental Setup and Analysis
Indic-TunedLens is trained using the Sarvam-1 model as the base model. All the implementations are carried out in Pytorch  [^2] https://pytorch.org/. To train and test Indic-TunedLens, we utilize two datasets that provide comprehensive multilingual coverage across Indian languages. We train our model on eleven languages, Bengali, English, Gujarati, Hindi, Kannada, Malayalam, Marathi, Panjabi, Nepali, Tamil and Telugu from the Sangraha dataset (Rahman Khan et al., 2024). For evaluation, we employ a curated subset of the multilingual Massive Multitask Language Understanding (MMLU) dataset (Dac Lai et al., 2023) adapted for Indian languages, comprising Bengali, English, Gujarati, Hindi, Kannada, Malayalam, Marathi, Nepali, Tamil and Telugu samples. The inclusion of Panjabi as a train only language enables assessment of cross lingual transfer capabilities to unseen but related languages. Detailed dataset statistics and implementation details are provided in Appendices A.1 and A.2, respectively.
@@
Figure 1: This figure shows the entropy heatmap of the standard Tuned Lens, which was developed for English-centric models. The high and irregular entropy across layers suggests unstable intermediate representations and weak alignment for Indian languages, with predictions biased toward English tokens.
Figure 2: Entropy heatmap for the Indic-TunedLens. Entropy decreases more smoothly across layers, indicating progressive information consolidation and improved semantic alignment, with intermediate predictions increasingly generating meaningful Hindi tokens.
Layer-wise Accuracy
We quantify alignment between intermediate representations and the final next-token prediction using a layer-wise agreement score. For a sequence of length $T$, the top-1 token predicted by a lens at layer $n$ for position $t$ is
$\hat{y}_{t}^{(n)}=\arg\max_{i\in V}\,p_{i}^{(n)}(\cdot\mid t).$ (4)
Agreement is measured against the base model’s prediction at the final layer $L$. This interpretability metric indicates how early sufficient information for the final decision emerges. Higher early-layer agreement reflects better alignment of intermediate states with the vocabulary, enabling faithful intermediate predictions.
@@ 5 Results
To assess the effectiveness of Indic-TunedLens for multilingual transformer interpretability, we conduct a comprehensive comparative analysis against the standard Logit Lens approach across 10 Indian languages (detailed in Appendix A).
@@
Layer-wise uncertainty reveals language-specific meaning formation. The standard Tuned Lens approach, originally designed for English centric models, demonstrates significant limitations when applied to Indian language processing, as illustrated in Figure 1. We show an example question (Figure 3) about monopoly concentration ratios from the Hindi MMLU dataset [^3] https://huggingface.co/datasets/alexandrainst/m_mmlu/ viewer/hi. The entropy (calculated as discussed in Section 4) heatmap which measures uncertainty in the model’s predicted token distribution at each layer reveals inconsistent patterns with high variance across layers, indicating poor alignment between the tuning mechanism and the model’s internal processing pathways for Indian languages. The visualization shows fragmented attention patterns, particularly in the middle layers (8-16), where semantic consolidation should occur, and scattered high entropy regions without clear progressive information refinement. This degraded performance stems from a projection mismatch: while the model internally encodes morphologically rich and semantically meaningful representations for Indian languages, the English-centric Tuned Lens fails to align these representations with the output vocabulary space.
@@
Maximum improvements occur in early layers (1-8) with language specific patterns reflecting morphological complexity. The heatmap in Figure 4 shows that Bengali, Nepali, and Hindi exhibit the strongest improvements in layers 1-4 (0.04-0.06 accuracy gain), corresponding to the morphological analysis phase where rich inflectional systems are processed. English shows concentrated improvements at layer 1 (0.06 gain) and more modest gains throughout middle layers, reflecting its analytic morphology. Gujarati and Kannada show very similar sustained improvements across layers 1-10 (0.02-0.04 gain), while Malayalam exhibits more distributed improvement patterns extending into middle layers (5-15). These layer specific improvement patterns enable linguistic hypotheses about where different linguistic phenomena are computed, suggesting that morphological analysis and early semantic composition occur in early layers, while deeper layers support more complex compositional and syntactic integration. This process happens earlier in languages like Telugu and Tamil, which form words by joining many meaningful parts together, compared to languages with more isolated word structures. The variation in improvement patterns across languages suggests that the learned transformations adapt to language specific processing dynamics, capturing the distinct computational trajectories by which different Indic languages resolve ambiguity and build compositional meaning. These findings demonstrate that interpretability methods cannot be language-agnostic: effective analysis of multilingual LLMs requires language-family-aware lenses that respect typological differences in morphological and compositional structure.
Indic-TunedLens demonstrates superior accuracy over Logit Lens across all Indian languages, with particularly pronounced improvements in early and middle layers. As shown in Figure 5, the standard Logit Lens exhibits minimal accuracy in early layers (0-15) across all languages, with values remaining below 0.02 for most Indian languages. The accuracy only begins to increase substantially after layer 20, reaching peak performance at layer 27 before dropping sharply at the final layer. In contrast, Indic-TunedLens shows immediate interpretability from layer 1, with English achieving 0.13 accuracy, followed by Telugu (0.11) and Gujarati (0.10). The framework maintains consistent accuracy between 0.03-0.06 throughout middle layers (5-20) for Indian languages, demonstrating robust intermediate representation capture. These findings suggest that language specific affine transformations enable meaningful interpretation of hidden states from the earliest layers of processing, whereas English centric approaches fail to capture the morphological richness and syntactic diversity inherent in Indian languages until much deeper in the network. The superior early layer performance of Indic-TunedLens indicates that specialized training on diverse Indic scripts creates representations better aligned with the model’s internal processing pathways for these languages, enabling more accurate decoding of intermediate semantic states. Individual language specific accuracy curves across all layers are provided in the appendix B.1.
@@
Indic-TunedLens consistently positions correct tokens at higher ranks across all layers, indicating improved confidence in predictions. Figure 6 reveals that Logit Lens maintains average ranks between 4-5 across layers 5-25 for most languages, with Malayalam showing particularly poor performance (ranks 4.5-5.8). The ranks only improve dramatically at layer 27, converging to 3.5-4 across languages. Conversely, Indic-TunedLens exhibits more favorable ranking patterns throughout the network, with most languages maintaining average ranks between 3.9-4.7 across middle layers. Notably, Hindi demonstrates superior rank positioning (4.5-5 in early layers, improving to 3.5-4.0 in later layers), while Malayalam shows the most substantial improvement from its Logit Lens baseline, achieving ranks of 4.2-4.8 in layers 10-25. The consistently better ranking performance of Indic-TunedLens across all languages demonstrates that the learned affine transformations more effectively align intermediate representations with the final vocabulary distribution, placing correct tokens among the top candidates earlier in the decoding process. This improved ranking behavior is particularly critical for morphologically rich languages where token level disambiguation requires integration of multiple linguistic features, and suggests that Indic-TunedLens better captures the hierarchical semantic refinement that occurs across transformer layers for Indian languages. Detailed rank distribution plots for each language are available in the appendix B.2.
Figure 7: Accuracy by Token Positions
@@ 7 Conclusion
Acknowledgments
Mamta gratefully acknowledges the support from the Engineering and Physical Sciences Research Council (EPSRC, grant number EP/X04162X/1).
Appendix A Experimental Setup
A.1 Dataset
@@ A.2 Implementation Details
Appendix B Detailed Analysis
B.1 Language Wise Layer-wise Accuracy Comparison between Indic-TunedLens and LogitLens
@@
Figure 17: Layer-wise Accuracy Comparison for Telugu
B.2 Language Wise Average Rank of Correct Predictions between Indic-TunedLens and LogitLens
@@
Figure 27: Average Rank of Correct Predictions for Telugu